3. Investment Horizons. How long does the Rule of 72 predict that it will take for $500 to double in value when the effective annual interest rate is 1 percent? Exactly how many years will it take for $500 to double in value when the effective annual interest rate is 1 percent?
Answer #1

1)

how long (using rule of 72)=72/rate of interest

=72/1

=72 years

2)

assume n=exactly number of years

future value=current value*(1+r)^n

(500*2)=500*(1+1%)^n

1.01^n=2

take natural log on both sides

n=ln(2)/ln(1.01)

=69.66 years

the above is answer..